OpenStudy (anonymous):

Let f(x)=5x^2. Find a value A such that the average rate of change of f(x) from 1 to A equals 60.

OpenStudy (amistre64):

the average rate of change is just (f(A) - f(1))/(A-1)

OpenStudy (anonymous):

i know!

OpenStudy (anonymous):

im getting the wrong answer for some reason though

OpenStudy (anonymous):

and im getting frustrated

OpenStudy (amistre64):

\[\frac{5a^2-5.1^2}{a-1}=60\] \[5a^2-5=60(a-1)\] \[5a^2-5=60a-60\] \[5a^2-60a-5+60=0\] \[5(a^2-120a+11)=0\]

OpenStudy (anonymous):

5.1?

OpenStudy (amistre64):

\[a=\frac{1}{2}(12\pm\sqrt{12^2-4(11)})\]

OpenStudy (amistre64):

its easier for me to get to the . then the *

myininaya (myininaya):

\[\frac{5a^2-5(1)^2}{a-1}=60=> \frac{5(a^2-1)}{a-1}=60\] \[=>\frac{5(a-1)(a+1)}{a-1}=60 => 5(a+1)=60=> a+1=12\]

OpenStudy (amistre64):

pfft!! thats cheating, and only gets one of them :)

OpenStudy (anonymous):

ITS 11!!!!!!!!!!!!!!! O MY GOD!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (amistre64):

144-44 = 100 sqrt100 = 10 12 +- 10 /2 /2 6 +- 5 = 1 and 11?

OpenStudy (amistre64):

cant be 1 so that leaves 11, yeah

myininaya (myininaya):

a cannot be 1 dear lol

myininaya (myininaya):

lol

OpenStudy (amistre64):

i was getting to that :)

myininaya (myininaya):

so my way gets all the possible values of A lol

OpenStudy (amistre64):

not the complex ones :P

myininaya (myininaya):

shhh lets not go there lol